Two point charges , q_{1} = 33.8nC and q_{2} = 12.6nC are located on the x-at x = 0 and x=1.00 m. respectively . Where on the x-axis the electric field equal to zero?


1
Expert's answer
2021-08-19T14:22:59-0400

Given:

"q_1=33.8*10^{-9}\:\rm C"

"q_2=12.6*10^{-9}\:\rm C"

"d=1.00\:\rm m"



The electric field at the point A is zero. Hence


"E_1=E_2"

"k\frac{q_1}{x^2}=k\frac{q_2}{(d-x)^2}"

"(d-x)\sqrt{q_1}=x\sqrt{q_2}"

"x=\frac{d\sqrt{q_1}}{\sqrt{q_1}+\sqrt{q_2}}=d*\frac{1}{1+\sqrt{q_2/q_1}}"

"x=1.00*\frac{1}{1+\sqrt{12.6/33.8}}=0.62\:\rm m"